# Towards a Unified Framework for Statistical and Mathematical Modeling

### v1

@@ Introduction
Figure 1: Visualization of types of identification results
@@
As indicated by this model, specific parameter values for $\theta$ and $\lambda$ must be chosen to compute $\psi$. Before considering how $\bar{\psi}(\theta,\lambda)$ changes for the input $\theta\in\Theta$ and $\lambda\in\Lambda$, some additional constraints are placed on the function $g$ and $h$. Here, the output of $g$ (i.e., image) covers the interval from zero to one for all unique values of $a$ (i.e., $\{g(a;\theta):\theta\in\Theta\}=[0,1]$ for $a\in\mathcal{A}$). Similarly, $\{h(a,m;\lambda):\lambda\in\Lambda\}=[0,1]$ for $a,m\in\mathcal{A}\times\mathcal{B}$. As examples of functions that satisfy these properties, consider $g(a;\theta):=\text{expit}(\theta_{0}+\theta_{1}a)$ with $\theta:=(\theta_{0},\theta_{1})\in\mathbb{R}^{2}$ and $h(a,m;\lambda):=\text{expit}(\lambda_{0}+\lambda_{1}a+\lambda_{2}m+\lambda_{3}am)$ with $\lambda:=(\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3})\in\mathbb{R}^{4}$, where $\text{expit}(x):=\{1+\exp(-x)\}^{-1}$. When adding the boundary points of zero and one, the set of outputs from the mathematical model will span the parameter space, i.e., $\{\bar{\psi}(\theta,\lambda):\theta,\lambda\in\Theta\times\Lambda\}\cup\{-1,0,1\}=[-1,1]=\Psi$, for any input $a,m$ combination. I refer to a mathematical model that satisfies this property as vacuous, since such a model provides no information about $\psi$ beyond that it lies within the parameter space for any value of $a$ and $m$ (which was known prior to constructing the model). Models that do not satisfy this property are referred to as non-vacuous. A vacuous model separates the implications of function choices from the parameter choices. Structural assumptions about the mechanism are thus encoded in the choice of parameters (e.g., $A$ not having a direct effect on $Y$ would be encoded by $\text{expit}(\lambda_{0}+0a+\lambda_{2}m+0am)$) and not by the choice of functions. Non-vacuous mathematical models instead rule out certain values for $\psi$ before even considering values of $\theta$ and $\lambda$.
@@
$\Pr_{S=1}(M^{a}=m\mid A=a)$ $\in\{g(a;\theta):\theta\in\Theta^{*}\}$ (2)
$\Pr_{S=1}(Y^{a}=1\mid M^{a}=m,A=a)$ $\in\{h(a,m;\lambda):\lambda\in\Lambda^{*}\}$
@@
which is a partial identification result for $\psi$.
@@
To identify and estimate $\psi$, a relative simple mathematical model is considered. Here, the following pharmacodynamic model is used
@@
$\bar{y}(b,a,m;\lambda):=b-\left[\lambda_{0}+a\lambda_{1}(\frac{m}{\lambda_{2}+m})+a\lambda_{3}\right],$
$I(\cdot)$ is the indicator function, $\lambda_{0}$ is a constant change in the baseline systolic blood pressure, $\lambda_{1}$ is the maximum response, $\lambda_{2}$ is the does with 50% of the maximal response, $m$ is the effective concentration, and $\lambda_{3}$ is the effect of amlodipine not through the effective concentration (e.g., changes in behavior due to side-effects).
@@
Having constructed and applied the model, now consider how the proposed framework can be used to judge this model and make recommendations for its improvement. A first concern revolves around the context. The scientific question addresses the place (United States) but not the time. Greater specificity of the scientific question is needed to assess whether the NHANES 2017-2018 data for $f(w)$ is reasonable. A second challenge is that the NHANES data only provides an estimate of the probability density function (albeit nonparametric). Within the mathematical model, the statistical uncertainty of NHANES being a sample is not incorporated into this mathematical model.
@@ Acknowledgements
Financial Support: The author of this work was supported by a grant from the National Institute of Allergy and Infectious Diseases (K01AI177102). The content is solely the responsibility of the author and does not necessarily represent the official views of the National Institutes of Health.

### v2

@@ Introduction
Figure 1: Visualization of different types of identification results
@@
As indicated by this model, specific parameter values for $\theta$ and $\lambda$ must be chosen to compute $\psi$. Before considering how $\bar{\psi}(\theta,\lambda)$ changes for the input $\theta\in\Theta$ and $\lambda\in\Lambda$, some additional constraints are placed on the function $g$ and $h$. Here, the output of $g$ (i.e., image) covers the interval from zero to one for all unique values of $a$ (i.e., $\{g(a;\theta):\theta\in\Theta\}=[0,1]$ for $a\in\mathcal{A}$). Similarly, $\{h(a,m;\lambda):\lambda\in\Lambda\}=[0,1]$ for $a,m\in\mathcal{A}\times\mathcal{B}$. As examples of functions that satisfy these properties, consider $g(a;\theta):=\text{expit}(\theta_{0}+\theta_{1}a)$ with $\theta:=(\theta_{0},\theta_{1})\in\mathbb{R}^{2}$ and $h(a,m;\lambda):=\text{expit}(\lambda_{0}+\lambda_{1}a+\lambda_{2}m+\lambda_{3}am)$ with $\lambda:=(\lambda_{0},\lambda_{1},\lambda_{2},\lambda_{3})\in\mathbb{R}^{4}$, where $\text{expit}(x):=\{1+\exp(-x)\}^{-1}$. When adding the boundary points of zero and one, the set of outputs from the mathematical model will span the parameter space, i.e., $\{\bar{\psi}(\theta,\lambda):\theta,\lambda\in\Theta\times\Lambda\}\cup\{-1,0,1\}=[-1,1]=\Psi$, for any input $a,m$ combination. I refer to a mathematical model that satisfies this property as vacuous, since such a model provides no information about $\psi$ beyond that it lies within the parameter space for any value of $a$ and $m$ (which was known prior to constructing the model). Models that do not satisfy this property are referred to as non-vacuous. A vacuous model separates the implications of function choices from the parameter choices. Structural assumptions about the mechanism are thus encoded in the choice of parameters (e.g., $A$ not having a direct effect on $Y$ would be encoded by $\text{expit}(\lambda_{0}+0a+\lambda_{2}m+0am)$) and not by the choice of functions. Non-vacuous mathematical models instead rule out certain values for $\psi$ before even considering values of $\theta$ and $\lambda$ (Figure 3).
Figure 3: Visualization of the parameter space covered by a vacuous versus non-vacuous model
@@
$\Pr_{S=1}(M^{a}=m)$ $\in\{g(a;\theta):\theta\in\Theta^{*}\}$ (2)
$\Pr_{S=1}(Y^{a}=1\mid M^{a}=m)$ $\in\{h(a,m;\lambda):\lambda\in\Lambda^{*}\}$
@@
which is a partial identification result for $\psi$. For intuition behind these varying results, see Figure 4
Figure 4: Visualization of the proposed types of identification results for mathematical models
@@
To identify and estimate $\psi$, a relative simple mathematical model is considered. Here, the following pharmacodynamic model based on the structure in Figure 5 is used
@@
$\bar{y}(b,a,m;\lambda):=b-\left[\lambda_{0}+a\lambda_{1}\left(\frac{m}{\lambda_{2}+m}\right)+a\lambda_{3}\right],$
$I(\cdot)$ is the indicator function, $\lambda_{0}$ is a constant change in the baseline systolic blood pressure, $\lambda_{1}$ is the maximum response, $\lambda_{2}$ is the dose with 50% of the maximal response, $m$ is the effective concentration, and $\lambda_{3}$ is the effect of amlodipine not through the effective concentration (e.g., changes in behavior due to side-effects).
Figure 5: Assumed mechanism between amlodipine and subsequent systolic blood pressure at 24 hours
@@
Having constructed and applied the model, now consider how the proposed framework can be used to judge this model and make recommendations for its improvement. A first concern revolves around the context. The scientific question addresses the place (United States) but not the time. Greater specificity of the scientific question is needed to assess whether the NHANES 2017-2018 data to parameterize $f(w)$ is reasonable. A second challenge is that the NHANES data only provides an estimate of the probability density function (albeit nonparametric). Within the mathematical model, the statistical uncertainty of NHANES being a sample is not incorporated into this mathematical model.
@@ Acknowledgements
Financial Support: The author of this work was supported by grants from the National Institute of Allergy and Infectious Diseases (K01AI177102) and the National Institute of General Medical Sciences (R01GM140564). The content is solely the responsibility of the author and does not necessarily represent the official views of the National Institutes of Health.